OpenStudy (anonymous):

0.6x +0.8 = 1.4 ? answer and how did you get it

OpenStudy (haichi):

-0.8 on both sides and u get 0.6x-0.6

OpenStudy (haichi):

0.6/0.6 on both sides u get x=1

OpenStudy (anonymous):

What is the answer nd how did you get it ?

OpenStudy (haichi):

the answer is 1

OpenStudy (haichi):

Read my explantion

OpenStudy (anonymous):

0.6x+0.8=1.4 0.6x=1.4-0.8 0.6x=0.6 x=0.6/0.6 x=1

OpenStudy (haichi):

0.6x+0.8=1.4--- _0.8 0n both sides and u get 0.6x=0.6

OpenStudy (haichi):

0.6/0.6 on both sides equals to 1

OpenStudy (anonymous):

anything / by its self = 1

OpenStudy (anonymous):

solve for x what you want to do is get x by itself on one side First Step subtract 0.8 from both sides because since u have+0.8 u do the opposite operation u get 0.6x+0.8-0.8=1.4-0.8 when you do the subtraction it will look like 0.6x=0.6 Second Step divide 0.6 on both sides 0.6x/0.6=0.6/0.6 then u have x=1